# Policy Optimization in the Linear Quadratic Gaussian Problem: A Frequency Domain Perspective

### v3

@@ 5.1 Optimization over the Space $\mathcal{RH}_{\infty}$
$\nabla J_{\mathcal{K}}(\mathbf{Q},Q)=2\left(\mathbf{S},\;\mathbf{e}_{(q+m_{1})\times(q+m_{2})}\left(\operatorname{Res}_{\Re(s)\leq 0}\bigl(\mathbf{S}\bigr),m_{1},m_{2}\right)\right),$
@@
$\times\mathbf{M}_{21}^{\mathcal{K}}(j\omega)\Big]\frac{d\omega}{2\pi}+2\operatorname{tr}(Res_{\Re(s)\leq 0}(\mathbf{S}(s))\Delta Q^{\top})+\mathcal{O}(\|(\Delta\mathbf{Q},\Delta Q)\|^{2}_{\mathbb{U}}).$
@@
$+2\operatorname{tr}(Res_{\Re(s)\leq 0}(\mathbf{S}(s))\Delta Q^{\top}).$
@@
$2\int_{-\infty}^{\infty}\operatorname{tr}\left[\Bigl(\mathbf{S}(j\omega)^{*}\Delta(\mathbf{Q}(j\omega))\right]\frac{d\omega}{2\pi}+2\operatorname{tr}(Res_{\Re(s)\leq 0}(\mathbf{S}(s))\Delta Q^{\top}).$
@@
$2\int_{-\infty}^{\infty}\operatorname{tr}\left[\Bigl(\mathbf{S}(j\omega)^{*}\Delta(\mathbf{Q}(j\omega))\right]\frac{d\omega}{2\pi}+2\operatorname{tr}(Res_{\Re(s)\leq 0}(\mathbf{S}(s))\Delta Q^{\top})$
$=$ $2\int_{-\infty}^{\infty}\operatorname{tr}\left[\Bigl(\mathbf{S}(j\omega)^{*}\Delta(\mathbf{Q}(j\omega))\right]\frac{d\omega}{2\pi}+2\operatorname{tr}(\mathbf{e}_{(q+m_{1})\times(q+m_{2})}\left(\operatorname{Res}_{\Re(s)\leq 0}\bigl(\mathbf{S}(s)\bigr),m_{1},m_{2}\right)\Delta Q^{\top}).$
@@
$\quad+2\operatorname{tr}\left(\operatorname{Res}_{\Re(s)\leq 0}\left((\mathbf{M}_{21}^{\mathcal{K}}(-s))^{\top}\mathbf{M}_{21}^{\mathcal{K}}(s)\Delta Q(\mathbf{M}_{12}^{\mathcal{K}}(-s))^{\top}\mathbf{M}_{12}^{\mathcal{K}}(s)\Delta Q^{\top}\right)\right).$
@@ 5.2 The Gradient-Based Method and Convergence Analysis
   Update $Q_{k+1}\leftarrow Q_{k}-\eta\,\mathbf{e}_{(q+m_{1})\times(q+m_{2})}(\mathrm{Res}_{\Re(s)\leq 0}(\mathbf{S}_{k}(s)),m_{1},m_{2})$;
@@
and the operator $\mathcal{F}$ is defined as follows. At the final iteration, let $G_{\mathcal{K}_{0}}$ denote the nominal closed-loop system (the generalized plant formed by interconnecting the original plant with the initial controller parameter $\mathcal{K}_{0}$). The update $\Delta\mathbf{K}_{N}$, designed to improve the performance, is then interconnected with $G_{\mathcal{K}_{0}}$ via a lower Linear Fractional Transformation (LFT), as illustrated in [42, Section 10.1]. The final controller $\mathcal{K}_{N}$ is then defined as the overall input-output map of the interconnected system, which can be expressed by
@@
Theorem 5.5 (Lipschitz Continuity of the Gradient).
The gradient of $J_{\mathcal{K}^{0}}$ is $L$-Lipschitz continuous on $\mathbb{U}$; that is, there exists a constant $L>0$ such that for any $(\mathbf{Q},Q),\,(\mathbf{Q}^{\prime},Q^{\prime})\in\mathbb{U}$,
@@
$\phi(\mathbf{Q},Q)$ $\leq\phi(\mathbf{Q}^{\prime},Q^{\prime})+\langle\nabla\phi(\mathbf{Q}^{\prime},Q^{\prime}),\,(\hat{\mathbf{Q}},\hat{Q})-(\mathbf{Q}^{\prime},Q^{\prime})\rangle_{\mathbb{U}}+\frac{L_{\mathcal{K}_{0}}}{2}\|(\hat{\mathbf{Q}},\hat{Q})-(\mathbf{Q}^{\prime},Q^{\prime})\|_{\mathbb{U}}^{2}.$
Taking $(\hat{\mathbf{Q}},\hat{Q})-(\mathbf{Q}^{\prime},Q^{\prime})=\nabla\phi(\mathbf{Q}^{\prime},Q^{\prime})$ into (70), we have
$\phi(\mathbf{Q},Q)\leq\phi(\mathbf{Q}^{\prime},Q^{\prime})-\frac{1}{2L_{\mathcal{K}_{0}}}\|\nabla\phi(\mathbf{Q}^{\prime},Q^{\prime})\|_{\mathbb{U}}^{2}.$ (70)
@@
$\phi(\mathbf{Q}^{\prime},Q^{\prime})\leq\phi(\mathbf{Q},Q)-\frac{1}{2L_{\mathcal{K}_{0}}}\|\nabla\phi(\mathbf{Q}^{\prime},Q^{\prime})\|_{\mathbb{U}}^{2}.$ (71)
@@
This result constitutes a nontrivial extension of the classic Lipschitz continuity (or smoothness) of the gradient of the cost functional in finite-dimensional Euclidean space (see, e.g., [20, Sec. 2.1]), to the infinite-dimensional setting, with two notable distinctions:
@@
The analysis in this section operated under the assumption that all the system parameters are known. We further show that Algorithm 1 is extensible to a data-driven setting. This is possible because the algorithm’s iterative steps do not require an explicit model of the system dynamics. Instead, the updates rely exclusively on a few key quantities—the transfer functions $\mathbf{M}_{22}^{\mathcal{K}_{0}}$ and $\mathbf{S}_{k}(s)$, and the residue $\mathbf{e}_{(q+m_{1})\times(q+m_{2})}(\mathrm{Res}_{\Re(s)\leq 0}(\mathbf{S}_{k}(s)),m_{1},m_{2})$ which can be estimated directly from input-output data. We provide the detailed procedures for this estimation in Appendix A and experimentally validate their effectiveness in subsection 6.2.
@@ 6.2 Example 2: Validation of Data-Driven Estimation Methods
This section validates the data-driven estimation methods proposed in Appendix  A, which constitute essential components of our algorithm. Specifically, we evaluate the accuracy of estimating: 1) the transfer function matrix $\mathbf{M}_{22}^{\mathcal{K}_{0}}(s)$, 2) the sensitivity function $\mathbf{S}_{k}(s)$, and 3) the residue term $\mathbf{e}_{(q+m_{1})\times(q+m_{2})}\left(\mathrm{Res}_{\Re(s)\leq 0}(\mathbf{S}_{k}(s)),m_{1},m_{2}\right)$ via zeroth-order methods.
@@ 6.2.4 Validation of Residue Estimation
Finally, we test the Monte Carlo based estimation of the residue term $\mathbf{e}_{3\times 3}\left(\mathrm{Res}_{\Re(s)\leq 0}(\mathbf{S}_{0}(s)),1,1\right)$ (from Section A.0.3). The key parameters for this simulation are set as:
@@ Appendix A Data-Driven Estimation Procedures for Key Operators
    the residue terms, defined as $\mathbf{e}_{(q+m_{1})\times(q+m_{2})}\left(\mathrm{Res}_{\Re(s)\leq 0}(\mathbf{S}_{k}(s)),m_{1},m_{2}\right).$
@@ A.0.3 Estimation of Residue Terms
To estimate the residue term $\mathbf{e}_{(q+m_{1})\times(q+m_{2})}\left(\mathrm{Res}_{\Re(s)\leq 0}(\mathbf{S}_{k}(s)),m_{1},m_{2}\right)$, we employ zeroth-order (gradient-free) methods [10, 21].
@@
$\nabla_{Q_{k}}J_{\mathcal{K}_{0}}(\mathbf{Q}_{k},Q_{k})=2\mathbf{e}_{(q+m_{1})\times(q+m_{2})}\left(\mathrm{Res}_{\Re(s)\leq 0}(\mathbf{S}_{k}(s)),m_{1},m_{2}\right).$ (86)
@@ Appendix B Second-Order Optimality Condition
$D_{\Delta}^{2}J\left(\widetilde{\mathcal{K}}(\lambda)\right)=0,\quad\forall\lambda\in\mathbb{R},\ \Delta\in\mathcal{M}_{p}.$ (89)
Moreover, if $\mathcal{K}$ is not a global optimal solution to problem (8), there exist $\Delta\in\mathcal{M}_{p}$ and $\lambda\in\mathbb{R}$ such that $D_{\Delta}^{2}J(\widetilde{\mathcal{K}}(\lambda))<0$.
@@
$D_{\Delta}^{2}J(\widetilde{\mathcal{K}}(\lambda))\equiv 0,\quad\forall\lambda\in\mathbb{R},\Delta\in\mathcal{M}_{p}$
@@
Theorem 2 of [41] implies that the first term $D_{\Delta^{(1)}+\Delta^{(2)}+\Delta^{(3)}}^{2}J\left(\widetilde{\mathcal{K}}(\lambda)\right)\equiv 0,\ \forall\lambda\in\mathbb{R},~\Delta\in\mathcal{M}_{p}$ if and only if $\mathbf{H}_{11}\equiv 0,\ \forall s\in\mathbb{C}$. We therefore only analyze the subsequent terms $D_{\Delta^{(4)}}^{2}J\left(\widetilde{\mathcal{K}}(\lambda)\right)$$D_{\Delta^{(1)}+\Delta^{(2)}}D_{\Delta^{(4)}}J\left(\widetilde{\mathcal{K}}(\lambda)\right)$ and $D_{\Delta^{(3)}}D_{\Delta^{(4)}}J\left(\widetilde{\mathcal{K}}(\lambda)\right)$. First, for $D_{\Delta^{(4)}}^{2}J\left(\widetilde{\mathcal{K}}(\lambda)\right)$, it holds that
@@
According to (92), it follows that $D_{\Delta^{(4)}}^{2}J\left(\widetilde{\mathcal{K}}(\lambda)\right)\equiv 0,\ \forall\lambda\in\mathbb{R},~\Delta\in\mathcal{M}_{p}$ if and only if $\mathbf{H}_{22}\equiv 0,\ \forall s\in\mathbb{C}$. By similar arguments, for $D_{\Delta^{(1)}+\Delta^{(2)}}D_{\Delta^{(4)}}J\left(\widetilde{\mathcal{K}}(\lambda)\right)$, we have
@@
Equation (93) implies that the condition $D_{\Delta^{(1)}+\Delta^{(2)}}D_{\Delta^{(4)}}J\left(\widetilde{\mathcal{K}}(\lambda)\right)\equiv 0,\ \forall\lambda\in\mathbb{R},~\Delta\in\mathcal{M}_{p}$ holds if and only if $\mathbf{H}_{12}\equiv\mathbf{H}_{21}\equiv 0,\ \forall s\in\mathbb{C}$. Furthermore, for $D_{\Delta^{(3)}}D_{\Delta^{(4)}}J\left(\widetilde{\mathcal{K}}(\lambda)\right)$, it holds that
@@
$+4\operatorname{tr}\left(\left[\begin{array}[]{cc}\Sigma_{\mathcal{K}}&0\\ 0&0\end{array}\right]P_{\mathcal{K},\Delta^{(3)}}^{\prime}\left[\begin{array}[]{ccc}0&0&0\\ 0&0&\Delta_{A1}\\ 0&\Delta_{A2}&0\end{array}\right]\right)\equiv 0,\ \forall\lambda\in\mathbb{R},\Delta\in\mathcal{M}_{p}.$
In conclusion, $D_{\Delta}^{2}J\left(\widetilde{\mathcal{K}}(\lambda)\right)\equiv 0$ for any $\lambda\in\mathbb{R},~\Delta\in\mathcal{M}_{p}$ if and only if the condition (27) holds. Meanwhile, based on the specific forms of the derivatives, when (89) does not hold, it is evident that there exist $\Delta\in\mathcal{M}_{p}$ and $\lambda\in\mathbb{R}$ such that $D_{\Delta}^{2}J(\widetilde{\mathcal{K}}(\lambda))<0$. This completes the proof. ∎

### v4

@@ 5.1 Optimization over the Space $\mathcal{RH}_{\infty}$
$\nabla J_{\mathcal{K}}(\mathbf{Q},Q)=2\left(\mathbf{S},\;\mathbf{e}_{(q+m_{1})\times(q+m_{2})}\left(\operatorname{Res}_{\Re(s)<0}\bigl(\mathbf{S}\bigr),m_{1},m_{2}\right)\right),$
@@
$\times\mathbf{M}_{21}^{\mathcal{K}}(j\omega)\Big]\frac{d\omega}{2\pi}+2\operatorname{tr}(\operatorname{Res}_{\Re(s)<0}(\mathbf{S}(s))\Delta Q^{\top})+\mathcal{O}(\|(\Delta\mathbf{Q},\Delta Q)\|^{2}_{\mathbb{U}}).$
@@
$+2\operatorname{tr}(\operatorname{Res}_{\Re(s)<0}(\mathbf{S}(s))\Delta Q^{\top}).$
@@
$2\int_{-\infty}^{\infty}\operatorname{tr}\left[\Bigl(\mathbf{S}(j\omega)^{*}\Delta(\mathbf{Q}(j\omega))\right]\frac{d\omega}{2\pi}+2\operatorname{tr}(\operatorname{Res}_{\Re(s)<0}(\mathbf{S}(s))\Delta Q^{\top}).$
@@
$2\int_{-\infty}^{\infty}\operatorname{tr}\left[\Bigl(\mathbf{S}(j\omega)^{*}\Delta(\mathbf{Q}(j\omega))\right]\frac{d\omega}{2\pi}+2\operatorname{tr}(\operatorname{Res}_{\Re(s)<0}(\mathbf{S}(s))\Delta Q^{\top})$
$=$ $2\int_{-\infty}^{\infty}\operatorname{tr}\left[\Bigl(\mathbf{S}(j\omega)^{*}\Delta(\mathbf{Q}(j\omega))\right]\frac{d\omega}{2\pi}+2\operatorname{tr}(\mathbf{e}_{(q+m_{1})\times(q+m_{2})}\left(\operatorname{Res}_{\Re(s)<0}\bigl(\mathbf{S}(s)\bigr),m_{1},m_{2}\right)\Delta Q^{\top}).$
@@
$\quad+2\operatorname{tr}\left(\operatorname{Res}_{\Re(s)<0}\left((\mathbf{M}_{21}^{\mathcal{K}}(-s))^{\top}\mathbf{M}_{21}^{\mathcal{K}}(s)\Delta Q(\mathbf{M}_{12}^{\mathcal{K}}(-s))^{\top}\mathbf{M}_{12}^{\mathcal{K}}(s)\Delta Q^{\top}\right)\right).$
@@ 5.2 The Gradient-Based Method and Convergence Analysis
   Update $Q_{k+1}\leftarrow Q_{k}-\eta\,\mathbf{e}_{(q+m_{1})\times(q+m_{2})}(\operatorname{Res}_{\Re(s)<0}(\mathbf{S}_{k}(s)),m_{1},m_{2})$;
@@
and the operator $\mathcal{F}$ is defined as follows. At the final iteration, let $G_{\mathcal{K}_{0}}$ denote the nominal closed-loop system (the generalized plant formed by interconnecting the original plant with the initial controller parameter $\mathcal{K}_{0}$). The update $\Delta\mathbf{K}_{N}$, designed to improve the performance, is then interconnected with $G_{\mathcal{K}_{0}}$ via a lower Linear Fractional Transformation (LFT), as illustrated in [44, Section 10.1]. The final controller $\mathcal{K}_{N}$ is then defined as the overall input-output map of the interconnected system, which can be expressed by
@@
Theorem 5.5.
There exists a constant $L>0$ such that for any $(\mathbf{Q},Q),\,(\mathbf{Q}^{\prime},Q^{\prime})\in\mathbb{U}$,
@@
$\phi({\mathbf{Q}},{Q})\leq\phi(\mathbf{Q}^{\prime},Q^{\prime})+\langle\nabla\phi(\mathbf{Q}^{\prime},Q^{\prime}),\,(\hat{\mathbf{Q}},\hat{Q})-(\mathbf{Q}^{\prime},Q^{\prime})\rangle_{\mathbb{U}}+\frac{L_{\mathcal{K}_{0}}}{2}\|(\hat{\mathbf{Q}},\hat{Q})-(\mathbf{Q}^{\prime},Q^{\prime})\|_{\mathbb{U}}^{2}.$ (72)
Taking $(\hat{\mathbf{Q}},\hat{Q})=\nabla\phi(\mathbf{Q}^{\prime},Q^{\prime})+(\mathbf{Q}^{\prime},Q^{\prime})$ into (72), we have
$\phi(\mathbf{Q},Q)\leq\phi(\mathbf{Q}^{\prime},Q^{\prime})-\frac{1}{2L_{\mathcal{K}_{0}}}\|\nabla\phi(\mathbf{Q}^{\prime},Q^{\prime})\|_{\mathbb{U}}^{2}$ (73)
which can be rewritten as
$J_{\mathcal{K}_{0}}(\mathbf{Q},Q)\leq J_{\mathcal{K}_{0}}(\mathbf{Q}^{\prime},Q^{\prime})-\left\langle\nabla J_{\mathcal{K}^{0}}(\mathbf{Q},Q),\,\Delta\right\rangle_{\mathbb{U}}-\frac{1}{2L_{\mathcal{K}_{0}}}\|\nabla\phi(\mathbf{Q}^{\prime},Q^{\prime})\|_{\mathbb{U}}^{2}$ (74)
@@
$J_{\mathcal{K}_{0}}(\mathbf{Q}^{\prime},Q^{\prime})\leq J_{\mathcal{K}_{0}}(\mathbf{Q},Q)+\left\langle\nabla J_{\mathcal{K}^{0}}(\mathbf{Q}^{\prime},Q^{\prime}),\,\Delta\right\rangle_{\mathbb{U}}-\frac{1}{2L_{\mathcal{K}_{0}}}\|\nabla\phi(\mathbf{Q}^{\prime},Q^{\prime})\|_{\mathbb{U}}^{2}$ (75)
@@
This result constitutes an extension of the classic Lipschitz continuity (or smoothness) of the gradient of the cost functional in finite-dimensional Euclidean space (see, e.g., [21, Sec. 2.1]), to the infinite-dimensional setting, with two notable distinctions:
@@
The analysis in this section operated under the assumption that all the system parameters are known. We further show that Algorithm 1 is extensible to a data-driven setting. This is possible because the algorithm’s iterative steps do not require an explicit model of the system dynamics. Instead, the updates rely exclusively on a few key quantities—the transfer functions $\mathbf{M}_{22}^{\mathcal{K}_{0}}$ and $\mathbf{S}_{k}(s)$, and the residue $\mathbf{e}_{(q+m_{1})\times(q+m_{2})}(\operatorname{Res}_{\Re(s)<0}(\mathbf{S}_{k}(s)),m_{1},m_{2})$ which can be estimated directly from input-output data. We provide the detailed procedures for this estimation in Appendix A and experimentally validate their effectiveness in subsection 6.2.
@@ 6.2 Example 2: Validation of Data-Driven Estimation Methods
This section validates the data-driven estimation methods proposed in Appendix  A, which constitute essential components of our algorithm. Specifically, we evaluate the accuracy of estimating: 1) the transfer function matrix $\mathbf{M}_{22}^{\mathcal{K}_{0}}(s)$, 2) the sensitivity function $\mathbf{S}_{k}(s)$, and 3) the residue term $\mathbf{e}_{(q+m_{1})\times(q+m_{2})}\left(\operatorname{Res}_{\Re(s)<0}(\mathbf{S}_{k}(s)),m_{1},m_{2}\right)$ via zeroth-order methods.
@@ 6.2.4 Validation of Residue Estimation
Finally, we test the Monte Carlo based estimation of the residue term $\mathbf{e}_{3\times 3}\left(\operatorname{Res}_{\Re(s)<0}(\mathbf{S}_{0}(s)),1,1\right)$ (from Section A.0.3). The key parameters for this simulation are set as:
@@ Appendix A Data-Driven Estimation Procedures for Key Operators
    the residue terms, defined as $\mathbf{e}_{(q+m_{1})\times(q+m_{2})}\left(\operatorname{Res}_{\Re(s)<0}(\mathbf{S}_{k}(s)),m_{1},m_{2}\right).$
@@ A.0.3 Estimation of Residue Terms
To estimate the residue term $\mathbf{e}_{(q+m_{1})\times(q+m_{2})}\left(\operatorname{Res}_{\Re(s)<0}(\mathbf{S}_{k}(s)),m_{1},m_{2}\right)$, we employ zeroth-order (gradient-free) methods [11, 22].
@@
$\nabla_{Q_{k}}J_{\mathcal{K}_{0}}(\mathbf{Q}_{k},Q_{k})=2\mathbf{e}_{(q+m_{1})\times(q+m_{2})}\left(\operatorname{Res}_{\Re(s)<0}(\mathbf{S}_{k}(s)),m_{1},m_{2}\right).$ (90)
@@ Appendix B Second-Order Optimality Condition
$D_{\Delta}^{2}J\left(\widetilde{\mathcal{K}}(\lambda)\right)=0,\quad\forall\lambda<0,\ \Delta\in\mathcal{M}_{p}.$ (93)
Moreover, if $\mathcal{K}$ is not a global optimal solution to problem (8), there exist $\Delta\in\mathcal{M}_{p}$ and $\lambda<0$ such that $D_{\Delta}^{2}J(\widetilde{\mathcal{K}}(\lambda))<0$.
@@
$D_{\Delta}^{2}J(\widetilde{\mathcal{K}}(\lambda))\equiv 0,\quad\forall\lambda<0,\Delta\in\mathcal{M}_{p}$
@@
Theorem 2 of [43] implies that the first term $D_{\Delta^{(1)}+\Delta^{(2)}+\Delta^{(3)}}^{2}J\left(\widetilde{\mathcal{K}}(\lambda)\right)\equiv 0,\ \forall\lambda<0,~\Delta\in\mathcal{M}_{p}$ if and only if $\mathbf{H}_{11}\equiv 0,\ \forall s\in\mathbb{C}$. We therefore only analyze the subsequent terms $D_{\Delta^{(4)}}^{2}J\left(\widetilde{\mathcal{K}}(\lambda)\right)$$D_{\Delta^{(1)}+\Delta^{(2)}}D_{\Delta^{(4)}}J\left(\widetilde{\mathcal{K}}(\lambda)\right)$ and $D_{\Delta^{(3)}}D_{\Delta^{(4)}}J\left(\widetilde{\mathcal{K}}(\lambda)\right)$. First, for $D_{\Delta^{(4)}}^{2}J\left(\widetilde{\mathcal{K}}(\lambda)\right)$, it holds that
@@
According to (96), it follows that $D_{\Delta^{(4)}}^{2}J\left(\widetilde{\mathcal{K}}(\lambda)\right)\equiv 0,\ \forall\lambda<0,~\Delta\in\mathcal{M}_{p}$ if and only if $\mathbf{H}_{22}\equiv 0,\ \forall s\in\mathbb{C}$. By similar arguments, for $D_{\Delta^{(1)}+\Delta^{(2)}}D_{\Delta^{(4)}}J\left(\widetilde{\mathcal{K}}(\lambda)\right)$, we have
@@
Equation (97) implies that the condition $D_{\Delta^{(1)}+\Delta^{(2)}}D_{\Delta^{(4)}}J\left(\widetilde{\mathcal{K}}(\lambda)\right)\equiv 0,\ \forall\lambda<0,~\Delta\in\mathcal{M}_{p}$ holds if and only if $\mathbf{H}_{12}\equiv\mathbf{H}_{21}\equiv 0,\ \forall s\in\mathbb{C}$. Furthermore, for $D_{\Delta^{(3)}}D_{\Delta^{(4)}}J\left(\widetilde{\mathcal{K}}(\lambda)\right)$, it holds that
@@
$+4\operatorname{tr}\left(\left[\begin{array}[]{cc}\Sigma_{\mathcal{K}}&0\\ 0&0\end{array}\right]P_{\mathcal{K},\Delta^{(3)}}^{\prime}\left[\begin{array}[]{ccc}0&0&0\\ 0&0&\Delta_{A1}\\ 0&\Delta_{A2}&0\end{array}\right]\right)\equiv 0,\ \forall\lambda<0,\Delta\in\mathcal{M}_{p}.$
In conclusion, $D_{\Delta}^{2}J\left(\widetilde{\mathcal{K}}(\lambda)\right)\equiv 0$ for any $\lambda<0,~\Delta\in\mathcal{M}_{p}$ if and only if the condition (27) holds. Meanwhile, based on the specific forms of the derivatives, when (93) does not hold, it is evident that there exist $\Delta\in\mathcal{M}_{p}$ and $\lambda<0$ such that $D_{\Delta}^{2}J(\widetilde{\mathcal{K}}(\lambda))<0$. This completes the proof. ∎